# NCL++: Nested Collaborative Learning for Long-Tailed Visual Recognition

### v1

@@ 3.2 Inter-Expert Collaborative Learning
To collaboratively learn multiple experts from each other, we employ the Inter-expert Collaborative Learning (InterCL) to allow each model to learn extra knowledge from others. We employ the knowledge distillation to allow the knowledge transferring among them, and following previous works zhang2018deep ; guo2020online , the Kullback Leibler (KL) divergence is employed to calculate the loss, which can be denoted as:
@@ 3.3 Intra-Expert Collaborative Learning
The learning uncertainty exists not only in different experts, but also in the same expert which takes different augmented images as the input. In other words, given an input image, if we augment it by several times and then feed them into the network, the corresponding outputs also vary greatly. Inspired by this, we further propose the Intra-Expert Collaborative Learning (IntraCL), which helps the network to alleviate the uncertainty within the expert. To be specific, for an input image ${\bf x}_{i}$, assume it will be augmented for $T$ times (like using RandAugment cubuk2020randaugment or other data augmentation strategies), and the augmented images are denoted as $\{{\bf x}_{i}^{t}\}_{t=1}^{T}$. All the augmented images are fed to the expert-$k$ and produce the corresponding balanced probabilities $\{{\bf p}({\bf x}_{i}^{t};{\bm{\theta}}_{k})\}_{t=1}^{T}$. Then, the knowledge distillation is performed to guide the network learning reliable and consistent outputs for different augmented copies. Moreover, the distillation also allows the knowledge transferring among the different augmented copies, which helps them to learn extra knowledge from each other. Similarly, KL divergence is adopted for loss calculation, and the corresponding loss of expert-$k$ can be represented as:
@@ 3.4 Hard Category Mining
In representation learning, one well-known and effective strategy to boost performance is Hard Example Mining (HEM) hermans2017defense . HEM claims that different examples have different importance, specifically, hard samples are more important than easy samples. Therefore, HEM selects hard samples for training, while discarding easy samples which contribute very little and are even detrimental to features learning. However, directly applying HEM to long-tailed visual recognition may distort the data distribution and make it more skewed in long-tailed learning. Although it can not be directly used for long-tailed learning, it still gives us an important inspiration, that is, do different categories have different degrees of importance in the training process?
@@ 4.1 Datasets and Protocols
We conduct experiments on five widely used datasets, including CIFAR10-LT cui2019class , CIFAR100-LT cui2019class , ImageNet-LT liu2019large , Places-LT zhou2017places , and iNaturalist 2018 van2018inaturalist .
CIFAR10-LT and CIFAR100-LT cui2019class are created from the original balanced CIFAR datasets krizhevsky2009learning . Specifically, the degree of data imbalance in datasets is controlled by the use of an Imbalance Factor (IF), which is defined by dividing the number of the most frequent category by that of the least frequent category. The imbalance factors of 100 and 50 are employed in these two datasets. ImageNet-LT liu2019large is sampled from the popular ImageNet dataset deng2009imagenet under long-tailed setting following the Pareto distribution with power value $\alpha$=6. ImageNet-LT contains 115.8K images from 1,000 categories. Places-LT is created from the large-scale dataset Places zhou2017places . This dataset contains 184.5K images from 365 categories. iNaturalist 2018 van2018inaturalist is the largest dataset for long-tailed visual recognition. iNaturalist 2018 contains 437.5K images from 8,142 categories, and it is extremely imbalanced with an imbalanced factor of 512.
According to previous works cui2019class ; kang2019decoupling the top-1 accuracy is employed for evaluation. Moreover, for iNaturalist 2018 dataset, we follow the works cai2021ace ; kang2019decoupling to divide classes into many (with more than 100 images), medium (with 20 $\sim$ 100 images) and few (with less than 20 images) splits, and further report the results on each split.
@@ 4.2 Implementation Details
For CIFAR10/100-LT, following cao2019learning ; zhang2021bag , we adopt ResNet-32 he2016deep as our backbone network and liner classifier for all the experiments. Input images are randomly cropped with size 32$\times$32. We utilize ResNet-50 he2016deep , ResNeXt-50 xie2017aggregated as our backbone network for ImageNet-LT, ResNet-50 for iNaturalist 2018 and pretrained ResNet-152 for Places-LT respectively, based on liu2019large ; kang2019decoupling ; cui2021parametric . Following zhang2021distribution , the cosine classifier is utilized for these models. Following previous works zhou2020bbn ; kang2019decoupling ; cui2021parametric , we resize the input image to 256$\times$256 pixels and take a 224$\times$224 crop from the original image or its horizontal flip. We use the SGD with a momentum of 0.9 and a weight decay of $2\times 10^{-4}$ as the optimizer to train all the models. As for experiments on CIFAR10/100-LT, the initial learning rate is 0.1 and decreases by 0.1 at epoch 320 and 360, respectively. The learning rate for Places-LT is 0.02 and decreases by 0.1 at epoch 10 and 20. For the rest datasets, the initial learning rate is set to 0.2 and decays by a cosine scheduler to $1\times 10^{-4}$. Unlike the previous works PaCo cui2021parametric and NCL li2022nested , we utilize less training epochs for ImageNet-LT and iNaturalist 2018, which is 200. And the training epochs for Places-LT is 30, same as previous works cui2021parametric ; kang2019decoupling . Cumulative gradient tan2022cross is utilized for ImageNet-LT and iNaturalist 2018. Specifically, the training batch size is set to 128, but the gradient keeps accumulating and the parameters are updated every two epochs. This trick uses less GPU memory to achieve similar training results as batch size of 256. Two experts are utilized for all datasets. Two augmented copies are utilized for ImageNet-LT and four for the rest. In addition, for fair comparison, following cui2021parametric , RandAugment cubuk2020randaugment is also used for all the experiments. The influence of RandAugment will be discussed in detail in the Sec. 4.4. These models are trained on 8 NVIDIA Tesla V100 GPUs. The $\beta=C_{hard}/C$ in HCM is set to 0.3, and the coefficient of InterCL and IntraCL loss $\lambda$ is set to 0.6. The influence of $\beta$ and $\lambda$ will be discussed in detail in the Sec. 4.4.
Table 1: Comparisons on CIFAR100-LT and CIFAR10-LT datasets with the IF of 100 and 50. ${}^{\dagger}$ indicates the ensemble performance is reported.
@@
CB Focal loss cui2019class CVPR’19 38.7 46.2 74.6 79.3
LDAM+DRW cao2019learning NeurIPS’19 42.0 45.1 77.0 79.3
LDAM+DAP jamal2020rethinking CVPR’20 44.1 49.2 80.0 82.2
BBN zhou2020bbn CVPR’20 39.4 47.0 79.8 82.2
LFME xiang2020learning ECCV’20 42.3
CAM zhang2021bag AAAI’21 47.8 51.7 80.0 83.6
Logit Adj. menon2020long ICLR’21 43.9 77.7
Xu et al. xu2021towards NeurIPS’21 45.5 51.1 82.8 84.3
LDAM+M2m kim2020m2m CVPR’21 43.5 79.1
MiSLAS zhong2021improving CVPR’21 47.0 52.3 82.1 85.7
LADE hong2021disentangling CVPR’21 45.4 50.5
Hybrid-SC wang2021contrastive CVPR’21 46.7 51.9 81.4 85.4
DiVE he2021distilling ICCV’21 45.4 51.3
SSD li2021self_iccv ICCV’21 46.0 50.5
PaCo cui2021parametric ICCV’21 52.0 56.0
xERM zhu2021cross AAAI’22 46.9 52.8
RISDA chen2021imagine AAAI’22 50.2 53.8 79.9 84.2
Batchformer hou2022batchformer CVPR’22 52.4
RIDE (4 ${}^{\dagger}$ wang2020long ICLR’21 49.1
ACE (4 ${}^{\dagger}$ cai2021ace ICCV’21 49.6 51.9 81.4 84.9
TLC (4 ${}^{\dagger}$ li2021trustworthy CVPR’22 49.8 80.4
NCL li2022nested CVPR’22 53.3 56.8 84.7 86.8
NCL (3 ${}^{\dagger}$ li2022nested CVPR’22 54.2 58.2 85.5 87.3
@@
NCL++ (2 ${}^{\dagger}$ 56.3 59.8 87.2 88.8
Table 2: Comparisons on ImageNet-LT and Places-LT datasets. ${}^{\dagger}$ indicates the ensemble performance is reported.
@@
OLTR liu2019large CVPR’19 34.1 35.9
BBN zhou2020bbn CVPR’20 48.3 49.3
NCM kang2019decoupling ICLR’20 35.5 44.3 47.3 36.4
cRT kang2019decoupling ICLR’20 41.8 47.3 49.6 36.7
$\tau$-norm kang2019decoupling ICLR’20 40.6 46.7 49.4 37.9
LWS kang2019decoupling ICLR’20 41.4 47.7 49.9 37.6
BSCE ren2020balanced NeurIPS’20 38.7
Xu et al. xu2021towards NeurIPS’21 42.9 48.4
DisAlign zhang2021distribution CVPR’21 52.9
DiVE he2021distilling ICCV’21 53.1
SSD li2021self_iccv ICCV’21 56.0
PaCo cui2021parametric ICCV’21 57.0 58.2 41.2
ALA Loss zhao2021adaptive AAAI’22 52.4 53.3 40.1
xERM zhu2021cross AAAI’22 54.1 39.3
RISDA chen2021imagine AAAI’22 50.7
MBJ liu2020memory AAAI’22 52.1 38.1
WD alshammari2022long CVPR’22 53.9
BatchFormer hou2022batchformer CVPR’22 47.6 57.4 41.6
RIDE (4 ${}^{\dagger}$ wang2020long ICLR’21 55.4 56.8
MBJ+RIDE (4 experts) liu2020memory AAAI’22 57.7
ACE (3 ${}^{\dagger}$ cai2021ace ICCV’21 44.0 54.7 56.6
TLC (4 ${}^{\dagger}$ li2021trustworthy CVPR’22 55.1
NCL li2022nested CVPR’22 46.8 57.4 58.4 41.5
NCL (3 ${}^{\dagger}$ li2022nested CVPR’22 47.7 59.5 60.5 41.8
@@
NCL++ (2 ${}^{\dagger}$ 49.9 59.6 60.9 42.4
Table 3: Comparisons on iNaturalist 2018 dataset with ResNet-50. ${}^{\dagger}$ indicates the ensemble performance is reported.
@@
OLTR liu2019large CVPR’19 59.0 64.1 64.9 63.9
BBN zhou2020bbn CVPR’20 49.4 70.8 65.3 66.3
DAP jamal2020rethinking CVPR’20 67.6
NCM kang2019decoupling ICLR’20
cRT kang2019decoupling ICLR’20 69.0 66.0 63.2 65.2
$\tau$-norm kang2019decoupling ICLR’20 65.6 65.3 65.9 65.6
LWS kang2019decoupling ICLR’20 65.0 66.3 65.5 65.9
LDAM+DRW cao2019learning NeurIPS’19 68.0
Logit Adj. menon2020long ICLR’21 66.4
CAM zhang2021bag AAAI’21 70.9
SSD li2021self_iccv ICCV’21
PaCo cui2021parametric ICCV’21 73.2
RIDE (3 ${}^{\dagger}$ wang2020long ICLR’21 70.9 72.4 73.1 72.6
ACE (3 ${}^{\dagger}$ cai2021ace ICCV’21 72.9
ALA Loss zhao2021adaptive AAAI’22 71.3 70.8 70.4 70.7
xERM zhu2021cross AAAI’22 67.3
RISDA chen2021imagine AAAI’22 69.1
MBJ liu2020memory AAAI’22 70.0
MBJ+RIDE liu2020memory AAAI’22 73.2
WD alshammari2022long CVPR’22 71.2 70.4 69.7 70.2
BatchFormer hou2022batchformer CVPR’22 65.5 74.5 75.8 74.1
NCL li2022nested CVPR’22 72.0 74.9 73.8 74.2
NCL (3 ${}^{\dagger}$ li2022nested CVPR’22 72.7 75.6 74.5 74.9
@@ 4.3 Comparisons to Prior Arts
We compare the proposed NCL++ with previous state-of-art methods, like BBN zhou2020bbn , RIDE wang2020long , NCL li2022nested . Besides, the baseline result of single network with using BSCE loss is also reported. Two experts is the default setting for NCL++, which has smaller model size than NCL, but achieves better performance. Comparisons on CIFAR10/100-LT are shown in Table 1, comparisons on ImageNet-LT and Places-LT are shown in Table 2, and comparisons on iNaturalist 2018 are shown in Table 3. Our proposed method achieves the state-of-the-art performance on all datasets. For only using a single expert for evaluation, our NCL++ outperforms previous methods on CIFAR10-LT, CIFAR100-LT, ImageNet-LT and Places-LT with accuracies of 86.1% (IF of 50), 54.8% (IF of 100), 58.0% (with ResNet-50) and 42.0%, respectively. Compared with NCL li2022nested , the proposed NCL++ achieves significant improvements on CIFAR-LT (54.8% vs. 53.3%) and ImageNet-LT with ResNet10 (49.0% vs. 46.8%). When further using an ensemble for evaluation, the performance on CIFAR10-LT, CIFAR100-LT, ImageNet-LT, Places-LT and iNaturalist2018 can be further improved to 87.2% (IF of 50), 56.3% (IF of 100), 59.6% (with ResNet-50), 42.4% and 75.2%, respectively. NCL++ uses fewer experts than NCL (2 experts vs. 3 experts), but achieves better ensemble performance on all the datasets. Similar to NCL, a single network can be used for evaluation, which will not bring extra computation but still outperforms previous multi-experts method, like ACE, TLC, etc.. When using ensemble of two experts, the performance has been further significantly improved.

### v3

@@ 3.2 Inter-Expert Collaborative Learning
To collaboratively learn multiple experts from each other, we employ the Inter-expert Collaborative Learning (InterCL) to allow each model to learn extra knowledge from others. We employ the knowledge distillation to allow the knowledge transferring among them, and following previous works [26, 24], the Kullback Leibler (KL) divergence is employed to calculate the loss, which can be denoted as:
@@ 3.3 Intra-Expert Collaborative Learning
The learning uncertainty exists not only in different experts, but also in the same expert which takes different augmented images as the input. In other words, given an input image, if we augment it by several times and then feed them into the network, the corresponding outputs also vary greatly. Inspired by this, we further propose the Intra-Expert Collaborative Learning (IntraCL), which helps the network to alleviate the uncertainty within the expert. To be specific, for an input image ${\bf x}_{i}$, assume it will be augmented for $T$ times (like using RandAugment [61] or other data augmentation strategies), and the augmented images are denoted as $\{{\bf x}_{i}^{t}\}_{t=1}^{T}$. All the augmented images are fed to the expert-$k$ and produce the corresponding balanced probabilities $\{{\bf p}({\bf x}_{i}^{t};{\bm{\theta}}_{k})\}_{t=1}^{T}$. Then, the knowledge distillation is performed to guide the network learning reliable and consistent outputs for different augmented copies. Moreover, the distillation also allows the knowledge transferring among the different augmented copies, which helps them to learn extra knowledge from each other. Similarly, KL divergence is adopted for loss calculation, and the corresponding loss of expert-$k$ can be represented as:
@@ 3.4 Hard Category Mining
In representation learning, one well-known and effective strategy to boost performance is Hard Example Mining (HEM) [62]. HEM claims that different examples have different importance, specifically, hard samples are more important than easy samples. Therefore, HEM selects hard samples for training, while discarding easy samples which contribute very little and are even detrimental to features learning. However, directly applying HEM to long-tailed visual recognition may distort the data distribution and make it more skewed in long-tailed learning. Although it can not be directly used for long-tailed learning, it still gives us an important inspiration, that is, do different categories have different degrees of importance in the training process?
@@ 4.1 Datasets and Protocols
We conduct experiments on five widely used datasets, including CIFAR10-LT [14], CIFAR100-LT [14], ImageNet-LT [11], Places-LT [9], and iNaturalist 2018 [63].
CIFAR10-LT and CIFAR100-LT [14] are created from the original balanced CIFAR datasets [64]. Specifically, the degree of data imbalance in datasets is controlled by the use of an Imbalance Factor (IF), which is defined by dividing the number of the most frequent category by that of the least frequent category. The imbalance factors of 100 and 50 are employed in these two datasets. ImageNet-LT [11] is sampled from the popular ImageNet dataset [7] under long-tailed setting following the Pareto distribution with power value $\alpha$=6. ImageNet-LT contains 115.8K images from 1,000 categories. Places-LT is created from the large-scale dataset Places [9]. This dataset contains 184.5K images from 365 categories. iNaturalist 2018 [63] is the largest dataset for long-tailed visual recognition. iNaturalist 2018 contains 437.5K images from 8,142 categories, and it is extremely imbalanced with an imbalanced factor of 512.
According to previous works [14, 17] the top-1 accuracy is employed for evaluation. Moreover, for iNaturalist 2018 dataset, we follow the works [22, 17] to divide classes into many (with more than 100 images), medium (with 20 $\sim$ 100 images) and few (with less than 20 images) splits, and further report the results on each split.
@@ 4.2 Implementation Details
For CIFAR10/100-LT, following [18, 30], we adopt ResNet-32 [1] as our backbone network and liner classifier for all the experiments. Input images are randomly cropped with size 32$\times$32. We utilize ResNet-50 [1], ResNeXt-50 [65] as our backbone network for ImageNet-LT, ResNet-50 for iNaturalist 2018 and pretrained ResNet-152 for Places-LT respectively, based on [11, 17, 66]. Following [41], the cosine classifier is utilized for these models. Following previous works [27, 17, 66], we resize the input image to 256$\times$256 pixels and take a 224$\times$224 crop from the original image or its horizontal flip. We use the SGD with a momentum of 0.9 and a weight decay of $2\times 10^{-4}$ as the optimizer to train all the models. As for experiments on CIFAR10/100-LT, the initial learning rate is 0.1 and decreases by 0.1 at epoch 320 and 360, respectively. The learning rate for Places-LT is 0.02 and decreases by 0.1 at epoch 10 and 20. For the rest datasets, the initial learning rate is set to 0.2 and decays by a cosine scheduler to $1\times 10^{-4}$. Unlike the previous works PaCo [66] and NCL [67], we utilize less training epochs for ImageNet-LT and iNaturalist 2018, which is 200. And the training epochs for Places-LT is 30, same as previous works [66, 17]. Cumulative gradient [68] is utilized for ImageNet-LT and iNaturalist 2018. Specifically, the training batch size is set to 128, but the gradient keeps accumulating and the parameters are updated every two epochs. This trick uses less GPU memory to achieve similar training results as batch size of 256. Two experts are utilized for all datasets. Two augmented copies are utilized for ImageNet-LT and four for the rest. In addition, for fair comparison, following [66], RandAugment [61] is also used for all the experiments. The influence of RandAugment will be discussed in detail in the Sec. 4.4. These models are trained on 8 NVIDIA Tesla V100 GPUs. The $\beta=C_{hard}/C$ in HCM is set to 0.3, and the coefficient of InterCL and IntraCL loss $\lambda$ is set to 0.6. The influence of $\beta$ and $\lambda$ will be discussed in detail in the Sec. 4.4.
Table 1: Comparisons on CIFAR100-LT and CIFAR10-LT datasets with the IF of 100 and 50. indicates the ensemble performance is reported.
@@
CB Focal loss [14] CVPR’19 38.7 46.2 74.6 79.3
LDAM+DRW [18] NeurIPS’19 42.0 45.1 77.0 79.3
LDAM+DAP [69] CVPR’20 44.1 49.2 80.0 82.2
BBN [27] CVPR’20 39.4 47.0 79.8 82.2
LFME [19] ECCV’20 42.3
CAM [30] AAAI’21 47.8 51.7 80.0 83.6
Logit Adj. [40] ICLR’21 43.9 77.7
Xu et al. [70] NeurIPS’21 45.5 51.1 82.8 84.3
LDAM+M2m [71] CVPR’21 43.5 79.1
MiSLAS [72] CVPR’21 47.0 52.3 82.1 85.7
LADE [39] CVPR’21 45.4 50.5
Hybrid-SC [73] CVPR’21 46.7 51.9 81.4 85.4
DiVE [74] ICCV’21 45.4 51.3
SSD [75] ICCV’21 46.0 50.5
PaCo [66] ICCV’21 52.0 56.0
xERM [76] AAAI’22 46.9 52.8
RISDA [77] AAAI’22 50.2 53.8 79.9 84.2
Batchformer [78] CVPR’22 52.4
RIDE (4 experts) [20] ICLR’21 49.1
ACE (4 experts) [22] ICCV’21 49.6 51.9 81.4 84.9
TLC (4 experts) [44] CVPR’22 49.8 80.4
NCL [67] CVPR’22 53.3 56.8 84.7 86.8
NCL (3 experts) [67] CVPR’22 54.2 58.2 85.5 87.3
@@
NCL++ (2 experts) 56.3 59.8 87.2 88.8
Table 2: Comparisons on ImageNet-LT and Places-LT datasets. indicates the ensemble performance is reported.
@@
OLTR [11] CVPR’19 34.1 35.9
BBN [27] CVPR’20 48.3 49.3
NCM [17] ICLR’20 35.5 44.3 47.3 36.4
cRT [17] ICLR’20 41.8 47.3 49.6 36.7
$\tau$-norm [17] ICLR’20 40.6 46.7 49.4 37.9
LWS [17] ICLR’20 41.4 47.7 49.9 37.6
BSCE [29] NeurIPS’20 38.7
Xu et al. [70] NeurIPS’21 42.9 48.4
DisAlign [41] CVPR’21 52.9
DiVE [74] ICCV’21 53.1
SSD [75] ICCV’21 56.0
PaCo [66] ICCV’21 57.0 58.2 41.2
ALA Loss [43] AAAI’22 52.4 53.3 40.1
xERM [76] AAAI’22 54.1 39.3
RISDA [77] AAAI’22 50.7
MBJ [79] AAAI’22 52.1 38.1
WD [80] CVPR’22 53.9
BatchFormer [78] CVPR’22 47.6 57.4 41.6
RIDE (4 experts) [20] ICLR’21 55.4 56.8
MBJ+RIDE (4 experts) [79] AAAI’22 57.7
ACE (3 experts) [22] ICCV’21 44.0 54.7 56.6
TLC (4 experts) [44] CVPR’22 55.1
NCL [67] CVPR’22 46.8 57.4 58.4 41.5
NCL (3 experts) [67] CVPR’22 47.7 59.5 60.5 41.8
@@
NCL++ (2 experts) 49.9 59.6 60.9 42.4
Table 3: Comparisons on iNaturalist 2018 dataset with ResNet-50. indicates the ensemble performance is reported.
@@
OLTR [11] CVPR’19 59.0 64.1 64.9 63.9
BBN [27] CVPR’20 49.4 70.8 65.3 66.3
DAP [69] CVPR’20 67.6
NCM [17] ICLR’20
cRT [17] ICLR’20 69.0 66.0 63.2 65.2
$\tau$-norm [17] ICLR’20 65.6 65.3 65.9 65.6
LWS [17] ICLR’20 65.0 66.3 65.5 65.9
LDAM+DRW [18] NeurIPS’19 68.0
Logit Adj. [40] ICLR’21 66.4
CAM [30] AAAI’21 70.9
SSD [75] ICCV’21
PaCo [66] ICCV’21 73.2
RIDE (3 experts) [20] ICLR’21 70.9 72.4 73.1 72.6
ACE (3 experts) [22] ICCV’21 72.9
ALA Loss [43] AAAI’22 71.3 70.8 70.4 70.7
xERM [76] AAAI’22 67.3
RISDA [77] AAAI’22 69.1
MBJ [79] AAAI’22 70.0
MBJ+RIDE [79] AAAI’22 73.2
WD [80] CVPR’22 71.2 70.4 69.7 70.2
BatchFormer [78] CVPR’22 65.5 74.5 75.8 74.1
NCL [67] CVPR’22 72.0 74.9 73.8 74.2
NCL (3 experts) [67] CVPR’22 72.7 75.6 74.5 74.9
@@ 4.3 Comparisons to Prior Arts
We compare the proposed NCL++ with previous state-of-art methods, like BBN [27], RIDE [20], NCL [67]. Besides, the baseline result of single network with using BSCE loss is also reported. Two experts is the default setting for NCL++, which has smaller model size than NCL, but achieves better performance. Comparisons on CIFAR10/100-LT are shown in Table 1, comparisons on ImageNet-LT and Places-LT are shown in Table 2, and comparisons on iNaturalist 2018 are shown in Table 3. Our proposed method achieves the state-of-the-art performance on all datasets. For only using a single expert for evaluation, our NCL++ outperforms previous methods on CIFAR10-LT, CIFAR100-LT, ImageNet-LT and Places-LT with accuracies of 86.1% (IF of 50), 54.8% (IF of 100), 58.0% (with ResNet-50) and 42.0%, respectively. Compared with NCL [67], the proposed NCL++ achieves significant improvements on CIFAR-LT (54.8% vs. 53.3%) and ImageNet-LT with ResNet10 (49.0% vs. 46.8%). When further using an ensemble for evaluation, the performance on CIFAR10-LT, CIFAR100-LT, ImageNet-LT, Places-LT and iNaturalist2018 can be further improved to 87.2% (IF of 50), 56.3% (IF of 100), 59.6% (with ResNet-50), 42.4% and 75.2%, respectively. NCL++ uses fewer experts than NCL (2 experts vs. 3 experts), but achieves better ensemble performance on all the datasets. Similar to NCL, a single network can be used for evaluation, which will not bring extra computation but still outperforms previous multi-experts method, like ACE, TLC, etc.. When using ensemble of two experts, the performance has been further significantly improved.